Six women and five men walk into an



appliance store at thesame time. There are five



salesperson available to help them.Find the



probability that a salesperson will first help at



least 2 women

1
Expert's answer
2022-02-27T16:41:52-0500

we define the total number of men and women = (5+6) = 11

We define (11C5)=462

let X be the number of women to be helped first

P((x≥2) = 1-(P(X=0)+P(X=1)

We define p(x=0)=( (6C0)*(5C5) ) /(11C5)=1/462

We define P(x=1)=( (6C1)*(5C4) ) /(11C5)= 30/462

Thus p(x≥2) = 1 - ( (1/462) + 30/462) ) = (431/462)=0.932900432